# Microcanonical simulated annealing: Massively parallel Monte Carlo simulations with sporadic random-number generation

### v2

@@ Abstract
I Introduction
@@
Optimization problems are indeed ubiquitous in everyday life —we all need to decide the most convenient order in which our daily tasks will be attended, or to determine how to allot our salary among the different chapters of a domestic budge— and Monte Carlo can be a precious method to address them. In a typical optimization problem, $N$ agents compete: think, for instance, of a traveling salesman that needs to visit $N$ cities exactly once, or, in the context of spin glasses, think of the assignment of the $N$ spins describing the magnetic configuration.
@@
Figure 1: After a time rescaling, our microcanonical simulated annealing reproduces canonical dynamics. We study the energy relaxation towards its equilibrium value by plotting the excess energy per spin over its equilibrium value, $e(t)-e_{\infty}$, versus simulation time $t$, as computed for a spin-glass sample on a $160\times 160\times 160$ lattice with binary couplings at temperature $T=0.7\approx 0.64T_{\mathrm{c}}$ ($T_{\mathrm{c}}$ is the critical temperature for the phase transition separating the paramagnetic phase at high temperatures from the spin-glass phase at low temperature). The sample is investigated using two dynamics: a standard, random-number greedy Metropolis algorithm (the label can in the plot stands for canonical; these data were obtained from the Janus II supercomputer) and the new microcanonical simulated annealing (MicSA) algorithm introduced in this work. Specifically, we are using the version of the algorithm with 6 walkers per spin, see Sec. II.8. We plot both data sets, subtracting the same estimate for $e_{\infty}$, which was obtained from the canonical simulation as explained in Appendix A. After the MicSA time is rescaled by the factor $k^{*}=6.034$ (see Table 1), $t^{\text{plot}}=k^{*}\times t^{\text{MicSA}}$, both energy relaxations are indistinguishable (the naif rescaling would have been $k^{\text{naif}}=6$, because of the 6 walkers per spin).
@@
In a canonical simulation, one may define an effective temperature (for instance) by comparing the instantaneous value of the energy with that of an equilibrated system at the effective temperature. Only in equilibrium does the effective temperature coincide with the thermal-bath temperature. In close analogy, see the more detailed discussion in Sect. II.7, the refresh of the daemons/walkers effectively cools the system only as long as thermal equilibrium has not been reached. In equilibrium, the net energy flow between spins and daemons/walkers cancels.
@@
The rescalability evinced in Fig. 1 can be related to analytical work on microcanonical dynamics —see, for instance Ref. [40] for model C critical dynamics (i.e., non-conserved order parameter and conserved energy). If the coherence length grows with time without bound —which is the case for a spin glass at $T<T_{\mathrm{c}}$, $T_{\mathrm{c}}$ being the critical temperature— and if the specific heat does not diverge in that limit (which is also the case for spin glasses), one expects that the canonical and the microcanonical dynamics can be superimposed at long times after a simple time rescaling. Although the algorithm proposed here is not microcanonical, it alternates between microcanonical and non-microcanonical phases. Hence, the field-theoretical prediction is reassuring.
@@
The structure of the rest of the paper is as follows: We recall the Edwards-Anderson model and explain our algorithm in Sec. II. Then, we compare the microcanonical and canonical dynamics for the Edwards-Anderson model in three dimensions in Sec. III, showing that both belong to the same dynamic universality class. We conclude by summarizing the results and discussing potential future applications of this algorithm in Sec. IV. We also provide additional information in two appendices. Appendix A provides technical details about our time-rescaling procedures, while Appendix B provides some basic information about our publicly available CUDA code.
II The model and the algorithm
@@
We have chosen to benchmark our algorithm in the rather demanding context of the three-dimensional Edwards-Anderson model [41, 42], which is the standard for spin glasses. While describing the model in Sec. II.1, we shall emphasize its separability, which is the crucial feature that makes parallel computation feasible. A similar separation between mutually non-interacting degrees of freedom is a necessary condition in any envisaged application of the solution proposed herein.
@@
We have chosen to organize this section in several separate paragraphs. After the somewhat introductory Secs. II.1 and II.2, each paragraph will discuss a specific decision we needed to make to design the algorithm. Indeed, efficient implementations require tuning Monte Carlo algorithms to the user’s specific needs, which are not only dictated by the peculiarities of the problem under study, but also by those of the hardware platform that will be used in the simulation. We hope that making our tuning process explicit will help the reader adapt the algorithm to their needs. We have summarized the defining details of the algorithms that we have numerically tested against a canonical simulation in Sec. II.8.
II.1 The model (and its separability)
@@
The crucial feature of the Hamiltonian $H^{\text{\tiny EA}}$ that makes parallel computations feasible is separability, which is itself a consequence of the limited interaction range (lattice nearest neighbors in our case). Here, separability means that the $N$ spins can be split into two classes such that spins in the same class are mutually non-interacting (entailing that they can be updated simultaneously). The number of classes could be larger than two, but it should be small to make the parallel computation efficient.
Now, let us address our specific problem. We note that the cubic lattice is bipartite: lattice sites can be classified as even or odd, depending on the parity of the sum of their Cartesian coordinates. Thus, with our nearest-neighbor interaction, odd sites interact solely with even sites (and vice versa). Hence, if we have fixed the values of the spins that occupy the (say) odd sublattice, the spins in the even sublattice are statistically independent random variables because the corresponding conditional probability extracted from the Boltzmann weight (2) factorizes. Furthermore, statistical independence carries out to the out-of-equilibrium context because in the (say) Metropolis algorithm, see below, the decision about flipping (or not) the spin at a site $\bm{x}$ is taken by looking only at the spins occupying its six lattice nearest-neighbor sites. The six neighbors have the opposite lattice parity; hence, if we keep fixed the spins in the (say) odd sublattice, the order in which we update the spins in the even sublattice is immaterial. In particular, the $N/2$ even spins can be updated simultaneously in a parallel simulation. With some effort, a domain decomposition similar to the odd and even sublattices can be found for essentially every short-ranged Hamiltonian. As we discuss in Sec. II.2, separability is the crucial feature endangered by the (in so many other ways fully satisfactory) Lustig microcanonical formalism.
II.1.1 Metropolis algorithm and detailed balance
Let us conclude this section by quickly reminding the reader of the Metropolis algorithm and the property of detailed balance it verifies for the canonical probability Eq. (2). The reader needing a more paused exposition may check, for instance, Refs. [53, 54].
For our purposes, Metropolis can be described as the following endless cycle composed of two steps (details may differ in particular implementations):
  1. 1.
    For every site $\bm{x}$ in the even sublattice:
    1. (a)
      Compute the change in $H^{\text{\tiny EA}}$, Eq. (1), if we flip the spin at site $\bm{x}$ (that is, $s_{\bm{x}}\to-s_{\bm{x}}$) while keeping all other spins fixed, and call it $\Delta H^{\text{\tiny EA}}$.
    2. (b)
      Draw a random number $R$ uniformly distributed in [0,1).
    3. (c)
      Flip the spin if
      $R<\text{min}\{1,\mathrm{e}^{-\beta\Delta H^{\text{\tiny EA}}}\}\,,$
      otherwise, leave the spin untouched.
  2. 2.
    Repeat procedure 1 for the odd sublattice.
The combination of steps 1 and 2 is often called a lattice sweep, because every spin of the lattice has had a chance to be flipped. Regarding step (1.b), note that in those cases where $\Delta H^{\text{\tiny EA}}\leq 0$ the spin will always get flipped. Hence, some physicists prefer to draw the random number $R$ only if $\Delta H^{\text{\tiny EA}}>0$ (which is reasonable in a serial update, but it is usually a bad idea in a parallel computation).
Detailed balance is a mathematical property that ensures that a particular probability density $\pi(X)$ is a stationary distribution for a particular Monte Carlo dynamics ($X$ is a point in configuration space, in our case, the whole spin configuration $X=\{s_{\bm{x}}\})$. Let ${\cal T}_{X\rightarrow Y}$ be the probability of going from configuration $X$ to configuration $Y$ in one step of a Monte Carlo algorithm (in the above Metropolis example, procedures 1 or 2 —half a lattice sweep– would be the basic step described by the matrix ${\cal T}$). Detailed balance states that,
$\pi(X)\,{\cal T}_{X\rightarrow Y}=\pi(Y)\,{\cal T}_{Y\rightarrow X}\,,$ (3)
for all pairs of states $X$ and $Y$. For instance, both procedures 1 and 2 of the Metropolis algorithm outlined above verify detailed balance when Boltzmann’s weight (2) plays the role of $\pi$. Two comments are in order: (i) detailed balance is a sufficient (but not necessary) condition for $\pi$ to be a stationary probability for dynamics ${\cal T}$  [^1] The really crucial feature is stationarity, $\sum_{X}\pi(X)\,{\cal T}_{X\rightarrow Y}=\pi(Y)$, which is implied by detailed balance. Actually, updating sequentially first the even and next the odd sites preserves stationarity but spoils detailed balance [53]. and (ii) to show that the stationary distribution for the dynamics ${\cal T}$ is unique, it is necessary to verify that an additional ergodicity condition named irreducibility is met [53].
II.2 A generalization of Lustig’s microcanonical ensemble
@@
We shall follow Lustig and expand the configuration space, introducing new degrees of freedom that we shall name hereafter either daemons or walkers (the distinction is discussed in Sect. II.6). Specifically, every spin will be endowed with a number $\gamma$ of these auxiliary degrees of freedom $n_{\bm{x},\alpha}\geq 0$, $\alpha=0,1,2,\ldots,\gamma-1$. The $n_{\bm{x},\alpha}$ could be either real or integer. Our total Hamiltonian will be [$H^{\text{\tiny EA}}$ was defined in Eq. (1)]
@@
where the Heaviside step function ensures that ${\cal E}>H^{\text{\tiny EA}}(\{s_{\bm{x}})\}$, and the modified partition function $\hat{Z}_{\text{mic}}$ ensures normalization. Eq. (8) is well adapted for a serial Monte Carlo simulation, where spins are updated one by one (an independent random number would be necessary per every spin update). Unfortunately, the effective action in Eq. (8), $\log({\cal E}-H_{\text{\tiny EA}})$ is highly non-linear in $H^{\text{\tiny EA}}$, which spoils the separability property discussed in Sec. II.1. Indeed, even an effective action quadratic in $H^{\text{\tiny EA}}$ would be non-separable because every pair of spins would be mutually interacting. With some ingenuity, a limited parallel computation can be set for an all-to-all interaction [7], but we regard this as a very undesirable feature. Hence, we shall depart from Lustig’s approach at this point.
II.3 Creutz’s move saves the day
@@
Of course, no random numbers are involved in the elementary move. However, it is amusing to notice that if a fresh value of the $n_{\bm{x},\alpha}$ were drawn for every spin-flip attempt  [^2] This is easy to do, although it would require a random number per spin update, because in the canonical ensemble, Eq. (6), the $n_{\bm{x},\alpha}$ are independent Poisson-distributed random variables., our elementary move would become the textbook Metropolis update —see, for instance, Sec. II.1.1 and Ref. [53]— for the spin system, Eq. (2). Baroque as it may seem, this disguising of the textbook Metropolis algorithm has resulted in a very efficient multispin coding implementation [57].
@@
II.4 Continuous or discrete daemons/walkers?
A computationally convenient feature of the $J=\pm 1$ choice in $H^{\text{\tiny EA}}$, Eq. (1), is that the energy changes that a single spin-flip may cause are limited. In three spatial dimensions, $\Delta H^{\text{\tiny EA}}=0,\pm 4,\pm 8$ or ±$\pm 12$. Hence, in our particular case, it is practical to limit the possible values of the daemons/walkers to integers multiple of four $n_{\bm{x},\alpha}=4m\,,\ m=0,1,2,3,\ldots$. For Gaussian couplings, it would probably be more natural to keep continuous daemons/walkers, or have them take values on a finer mesh of energies.
@@ II.4 Continuous or discrete daemons/walkers?
II.5 Daemons/walkers with a limited capacity
To further parallelize our computation we employ multispin coding [58]. We use the different bits in a memory word to represent different physical spins. Hence, we have chosen to restrict the possible values to $n_{\bm{x},\alpha}=4m$ with $m=0,1,2,3$. Thus $m$ is representable with two bits only, and we can use two 32-bit words to represent 32 different daemons/walkers.
@@ II.6 Daemons or walkers?
Interestingly enough, Refs. [34, 35] implemented their own daemon wandering (although they refreshed their daemons far more often than we do, see Sect. II.7).
II.7 The microcanonical simulated annealing
The operations described so far do not allow us to approach the equilibrium canonical distribution, Eq. (6). Indeed, with our microcanonical update, the total energy ${\cal E}$ remains equal to its starting value ${\cal E}_{\text{start}}$. Eq. (4) tells us that ${\cal E}_{\text{start}}$ is the sum of two contributions, the starting value of the spin Hamiltonian $H^{\text{\tiny EA}}$ (1) —typically much larger than the equilibrium value at inverse temperature $\beta,\langle H^{\text{\tiny EA}}\rangle_{\beta}$, and the energy of the daemons/walkers. Although the daemons/walkers will be directly initialized in thermal equilibrium through Eq. (12), it is clear that ${\cal E}_{\text{start}}$ will be too high. As a result, the microcanonical updates will lower the energy of the spins—which is desirable— at the cost of heating the daemons/walkers —which is undesirable. It is easy to assign an instantaneous temperature to the daemons/walkers, because the $\beta$-dependent equilibrium expectation value can be easily computed from Eq. (12). We have $\gamma$ daemons/walkers per each of the $N$ spins, as
@@ II.7 The microcanonical simulated annealing
Now, if we aim to reach the canonical equilibrium, Eq. (6), it is clear that we need to find a way to cool the daemons/walkers. In some ways, our solution will be similar to the strategy known as simulated annealing [17]. At specified times, see the discussion below, we throw away the daemon/walker configuration —leaving spins untouched— and generate a new, statistically independent one from Eq. (12). By construction, the new daemons/walkers configuration will have the correct effective temperature dictated by Eq. (13). In other words, the nominal daemons/walkers temperature $T$ will coincide with the effective temperature. Hence, we shall have decreased the total energy ${\cal E}$ of our annealing system. When equilibrium is finally reached, the energy gains and losses suffered by the spins finally compensate, and, although the effective temperature of the daemons/walkers fluctuates, it remains stable (on average).
From the above discussion, it should be clear that refreshing the daemons/walkers configuration at regular times, say every 1000 timesteps (see Sec. II.8 for the precise definition of timestep), should eventually lead any system of finite size to canonical equilibrium. This would dilute the need for random numbers by a factor of 1000, which would probably make the computational weight of generating random numbers irrelevant for most (if not all) applications. However, we shall try to be even thriftier with our random numbers. Indeed, the energy relaxation of spin glasses at and below their critical temperature decreases very slowly with time $t$. The excess energy over the equilibrium value decays as a power law, $H^{\text{\tiny EA}}(t)-\langle H^{\text{\tiny EA}}\rangle_{\beta}\propto N/t^{\epsilon(T)}$, where $\epsilon(T)$ is a fairly small exponent [59, 47] (the exponent is so small that the power law resembles a logarithmic behavior). In fact, for a law $1/\log(t)$, the time required to have a prefixed energy factor decay $f$ grows exponentially with $1/f$. We decided accordingly to refresh our daemons/walkers with At each of those timesteps, the spin system will experience an abrupt, but tiny, energy change as a reaction to the cooler daemons/walkers. This is inappreciable at the scale of Fig. 1, but can be observed with the more refined analysis in Fig. 2-d.
II.8 Putting all pieces together: the algorithms we tested

### v3

@@ Abstract
journal: Computer Physics Communications
1 Introduction
@@
Optimization problems are indeed ubiquitous in everyday life —we all need to decide the most convenient order in which our daily tasks will be attended, or to determine how to allot our salary among the different chapters of a domestic budget — and Monte Carlo can be a precious method to address them. In a typical optimization problem, $N$ agents compete: think, for instance, of a traveling salesman that needs to visit $N$ cities exactly once, or, in the context of spin glasses, think of the assignment of the $N$ spins describing the magnetic configuration.
@@
Figure 1: After a time rescaling, our microcanonical simulated annealing reproduces canonical dynamics. We study the energy relaxation towards its equilibrium value by plotting the excess energy per spin over its equilibrium value, $e(t)-e_{\infty}$, versus simulation time $t$, as computed for a spin-glass sample on a $160\times 160\times 160$ lattice with binary couplings at temperature $T=0.7\approx 0.64T_{\mathrm{c}}$ ($T_{\mathrm{c}}$ is the critical temperature for the phase transition separating the paramagnetic phase at high temperatures from the spin-glass phase at low temperature). The sample is investigated using two dynamics: a standard, random-number greedy Metropolis algorithm (the label can in the plot stands for canonical; these data were obtained from the Janus II supercomputer) and the new microcanonical simulated annealing (MicSA) algorithm introduced in this work. Specifically, we are using the version of the algorithm with 6 walkers per spin, see Sec. 2.8. We plot both data sets, subtracting the same estimate for $e_{\infty}$, which was obtained from the canonical simulation as explained in  A. After the MicSA time is rescaled by the factor $k^{*}=6.034$ (see Table 1), $t^{\text{plot}}=k^{*}\times t^{\text{MicSA}}$, both energy relaxations are indistinguishable (the naif rescaling would have been $k^{\text{naif}}=6$, because of the 6 walkers per spin).
@@
In a canonical simulation, one may define an effective temperature (for instance) by comparing the instantaneous value of the energy with that of an equilibrated system at the effective temperature. Only in equilibrium does the effective temperature coincide with the thermal-bath temperature. In close analogy, see the more detailed discussion in Sect. 2.7, the refresh of the daemons/walkers effectively cools the system only as long as thermal equilibrium has not been reached. In equilibrium, the net energy flow between spins and daemons/walkers cancels.
@@
The rescalability evinced in Fig. 1 can be related to analytical work on microcanonical dynamics —see, for instance, the Chapter entitled “Critical Dynamics” in Ref. [40] for model-C critical dynamics (i.e., non-conserved order parameter and conserved energy). If the coherence length grows with time without bound —which is the case for a spin glass at $T<T_{\mathrm{c}}$, $T_{\mathrm{c}}$ being the critical temperature— and if the specific heat does not diverge in that limit (which is also the case for spin glasses), one expects that the canonical and the microcanonical dynamics can be superimposed at long times after a simple time rescaling. Although the algorithm proposed here is not microcanonical, it alternates between microcanonical and non-microcanonical phases. Hence, the field-theoretical prediction is reassuring.
@@
The structure of the rest of the paper is as follows: We recall the Edwards-Anderson model and explain our algorithm in Sec. 2. Then, we compare the microcanonical and canonical dynamics for the Edwards-Anderson model in three dimensions in Sec. 3, showing that both belong to the same dynamic universality class. We conclude by summarizing the results and discussing potential future applications of this algorithm in Sec. 4. We also provide additional information in three appendices. A provides technical details about our time-rescaling procedures,  B provides some basic information about our publicly available CUDA code. Finally, C briefly describes the application of the method in equilibrium simulations for the specific example of a small sample of a three-dimensional spin glass with Gaussian couplings at low temperatures.
2 The model and the algorithm
@@
We have chosen to benchmark our algorithm in the rather demanding context of the three-dimensional Edwards-Anderson model [41, 42], which is the standard for spin glasses. While describing the model in Sec. 2.1, we shall emphasize its separability, which is the crucial feature that makes parallel computation feasible. A similar separation between mutually non-interacting degrees of freedom is a necessary condition in any envisaged application of the solution proposed herein.
@@
We have chosen to organize this section in several separate paragraphs. After the somewhat introductory Secs. 2.1 and 2.2, each paragraph will discuss a specific decision we needed to make to design the algorithm. Indeed, efficient implementations require tuning Monte Carlo algorithms to the user’s specific needs, which are not only dictated by the peculiarities of the problem under study, but also by those of the hardware platform that will be used in the simulation. We hope that making our tuning process explicit will help the reader adapt the algorithm to their needs. We have summarized the defining details of the algorithms that we have numerically tested against a canonical simulation in Sec. 2.8.
2.1 The model (and its separability)
@@
A crucial feature of the Hamiltonian $H^{\text{\tiny EA}}$ that makes parallel computations feasible is separability, which is itself a consequence of the limited interaction range (lattice nearest neighbors in our case). Here, separability means that the $N$ spins can be split into disjoint sets such that spins in the same set are mutually non-interacting (entailing that they can be updated simultaneously). The number of non-interacting sets could be larger than two, but it should be small to make parallel computing efficient. In our case, a very natural possibility is to group the spins into even and odd sublattices, depending on the parity of the sum of their Cartesian coordinates. This choice results in a checkerboard decomposition of the lattice which has proven instrumental in many computational and theoretical studies for decades (Ref. [55] provides an early example).
As we discuss in Sec. 2.2, separability is endangered by the (in so many other ways fully satisfactory) Lustig microcanonical formalism.
2.2 A generalization of Lustig’s microcanonical ensemble
@@
Initially, we shall follow Lustig and expand the configuration space, introducing new degrees of freedom that we shall name hereafter either daemons or walkers (the distinction is discussed in Sect. 2.6). Specifically, every spin will be endowed with a number $\gamma$ of these auxiliary degrees of freedom $n_{\bm{x},\alpha}\geq 0$, $\alpha=0,1,2,\ldots,\gamma-1$. The $n_{\bm{x},\alpha}$ could be either real or integer. Our total Hamiltonian will be $H^{\text{\tiny EA}}$ was defined in Eq. (1)]
@@
where the Heaviside step function ensures that ${\cal E}>H^{\text{\tiny EA}}(\{s_{\bm{x}})\}$, and the modified partition function $\hat{Z}_{\text{mic}}$ ensures normalization. Eq. (7) is well adapted for a serial Monte Carlo simulation, where spins are updated one by one (an independent random number would be necessary per every spin update). Unfortunately, the effective action in Eq. (7), $\log({\cal E}-H_{\text{\tiny EA}})$ is highly non-linear in $H^{\text{\tiny EA}}$, which spoils the separability property discussed in Sec. 2.1. Indeed, even an effective action quadratic in $H^{\text{\tiny EA}}$ would be non-separable because every pair of spins would be mutually interacting. With some ingenuity, a limited parallel computation can be set for an all-to-all interaction [7], but we regard this as a very undesirable feature. Hence, we shall depart from Lustig’s approach at this point.
2.3 Creutz’s move saves the day
@@
Of course, no random numbers are involved in the elementary move. However, it is amusing to notice that if a fresh value of the $n_{\bm{x},\alpha}$ were drawn for every spin-flip attempt, [^1] This is easy to do, although it would require a random number per spin update, because in the canonical ensemble, Eq. (5), the $n_{\bm{x},\alpha}$ are independent exponentially-distributed random variables. our elementary move would become the textbook Metropolis update —see, e.g., Refs. [53, 54]— for the spin system, recall Eq. (2). Baroque as it may seem, this disguising of the textbook Metropolis algorithm has resulted in a very efficient multispin coding implementation [56].
@@
2.4 Continuous or discrete daemons/walkers?
A computationally convenient feature of the $J=\pm 1$ choice in $H^{\text{\tiny EA}}$, Eq. (1), is that the energy changes that a single spin-flip may cause are limited. In three spatial dimensions, $\Delta H^{\text{\tiny EA}}=0,\pm 4,\pm 8$ or $\pm 12$. Hence, in our particular case, it is practical to limit the possible values of the daemons/walkers to integers multiple of four $n_{\bm{x},\alpha}=4m\,,\ m=0,1,2,3,\ldots$. For Gaussian couplings, it would probably be more natural to keep continuous daemons/walkers, or have them take values on a finer mesh of energies.
@@ 2.4 Continuous or discrete daemons/walkers?
2.5 Daemons/walkers with a limited capacity
To further parallelize our computation we employ multispin coding [57]. We use the different bits in a memory word to represent different physical spins. Hence, we have chosen to restrict the possible values to $n_{\bm{x},\alpha}=4m$ with $m=0,1,2,3$. Thus $m$ is representable with two bits only, and we can use two 32-bit words to represent 32 different daemons/walkers. Note that these choices are problem specific. In other situations, different choices are possible, see, for instance, C.
@@ 2.6 Daemons or walkers?
Interestingly enough, Refs. [34, 35] implemented their own daemon wandering (although they refreshed their daemons far more often than we do, see Sect. 2.7).
2.7 The microcanonical simulated annealing
[…]
@@ 2.7 The microcanonical simulated annealing
[…]
From the above discussion, it should be clear that refreshing the daemons/walkers configuration at regular times, say every 1000 timesteps (see Sec. 2.8 for the precise definition of timestep), should eventually lead any system of finite size to canonical equilibrium. This would dilute the need for random numbers by a factor of 1000, which would probably make the computational weight of generating random numbers irrelevant for most (if not all) applications. However, we shall try to be even thriftier with our random numbers. Indeed, the energy relaxation of spin glasses at and below their critical temperature decreases very slowly with time $t$. The excess energy over the equilibrium value decays as a power law, $H^{\text{\tiny EA}}(t)-\langle H^{\text{\tiny EA}}\rangle_{\beta}\propto N/t^{\epsilon(T)}$, where $\epsilon(T)$ is a fairly small exponent [58, 47] (the exponent is so small that the power law resembles a logarithmic behavior). In fact, for a law $1/\log(t)$, the time required to have a prefixed energy factor decay $f$ grows exponentially with $1/f$. We decided accordingly to refresh our daemons/walkers with At each of those timesteps, the spin system will experience an abrupt, but tiny, energy change as a reaction to the cooler daemons/walkers. This is inappreciable at the scale of Fig. 1, but can be observed with the more refined analysis in Fig. 2-d.
2.8 Putting all pieces together: the algorithms we tested